# Multiscale Modeling of Vacancy-Cluster Interactions and Solute Clustering Kinetics in Al-Mg-Zn Alloys

### v1

@@ Results
where $E^{\text{site}}_{\text{bind},i}$ is the energy of vacancy on site $i$. The elements, $\tau_{i}$, in the vector $\vec{\tau}$ represent the expected escape one-step time for the particular site. For each site $i$, $\tau_{i}$ accounts for all possible transitions to neighboring states. Migration barriers $E_{a{ij}}$, and attempt frequencies $\nu$ are incorporated to estimate escape times from each transient state.
$\tau_{i}=\left(\nu\sum_{j}{\exp\left(-\frac{E^{\text{bond}}_{\text{mig},ij}}{k_{B}T}\right)}\right)^{-1}.$ (5)
Here $\nu$ is set at $10^{13}$ Hz, as a typical atomic vibration frequency value in a crystal.
@@
Figure 4: KMC simulations for the verification of the absorbing Markov chain model. (a) The strategy for applying the KMC algorithm to determine the escaped state of a vacancy from a solute cluster. Open red circles denote vacancies. Blue, green, and orange circles represent Al, Mg, and Zn atoms, respectively. White stars mark the first nearest neighbor shell of the cluster. (b) The results comparison of the escape time $t_{\text{esc}}$ defined in Eq. 2 from the absorbing Markov chain model and from KMC verification for solute clusters generated by SA with different cluster sizes $N^{\text{clu}}$.
@@
These energetics applied in the absorbing Markov chain model, $E_{\text{bind}}^{\text{site}}$ and $E_{\text{mig}}^{\text{bond}}$, provide sufficient information to construct an accurate PEL for single-component clusters, such as Cu clusters in Fe-Cu systems, or Cu Guinier-Preston (GP) zones in Al-Cu alloys. In such systems, vacancy migration inside clusters does not induce structural rearrangements due to the indistinguishable nature of solute atoms. However, for binary clusters in multicomponent systems, such as Mg-Zn clusters in Al-Mg-Zn alloys, we recognize that local atomic rearrangements by vacancy migration may complicate the actual PEL. Despite these structural changes, we make a simplified-PEL assumption about the PEL. Specifically, we assume that the PEL experienced by a migrating vacancy can be approximated by the static ground-state structures of the cluster (obtained by SA). This simplification is reasonable at lower temperatures, where thermodynamics favors highly ordered, low-energy structures. By making this assumption, we effectively reduce the complexity of the PEL, enabling a computationally efficient absorbing Markov chain model to describe vacancy dynamics inside the clusters. This simplification maintains analytical tractability and computational efficiency, effectively capturing essential vacancy behavior without explicitly modeling structural rearrangements.
@@
To validate the absorbing Markov chain model and test the assumption of a simplified PEL, we applied KMC simulations. KMC and Markov chains share the same memorylessness property, which assumes that the next state depends only on the current state and not on past states. KMC calculates transition rates and selects events probabilistically, updating the system state and time step by step. Unlike the Markov model, KMC inherently accounts for structural rearrangements and complex PELs by simulating vacancy migration.
@@
Fig. 4 presents the verification for the clusters generated and optimized from the SA method (see Supplementary Note 3 for the clusters extracted from KMC simulations). Each cluster contains one vacancy and maintains a fixed Zn/Mg ratio of 2. Then, we performed KMC simulations on these clusters, focusing on determining the real escape times required for a vacancy to leave the cluster. In KMC simulations, the vacancies are tracked as they migrate through the cluster until they reach the surface, where they are considered to have escaped. This stopping criterion, illustrated in Fig. 4a, mirrors the definition used in the absorbing Markov chain model, providing a consistent basis for comparison between the two approaches. The KMC simulations were conducted at an elevated temperature of 450 K. This higher temperature facilitates fast transitions, enabling vacancies to overcome energy barriers and escape the energy basin more efficiently compared to simulations performed at room temperature. This higher temperature ensures that statistically meaningful results can be obtained within a feasible computational time. Fig. 4b compares the escape times predicted by the absorbing Markov chain model (Eq. 2) with those obtained from the KMC simulations. Here, results from absorbing Markov chain and KMC validation exhibit strong agreement, demonstrating that the simplified PEL approximation employed in the absorbing Markov model effectively captures the essential features of vacancy dynamics.
@@
Figure 5: Analysis for vacancy trapping ability of solute clusters using the Markov chain model. (a) Estimated vacancy escape time $t_{\text{esc}}$ at 300 K of solute clusters in terms of cluster size $N^{\text{clu}}$. (b) A conceptual figure to explain the singularity approximation for an energy well generated by a solute cluster. (c) Effective vacancy binding energy of a cluster $\tilde{E}^{\text{clu}}_{\text{bind}}$ defined in Eq. 7 in terms of cluster size $N^{\text{clu}}$.
@@
Accordingly, the estimated vacancy escape times at 300 K, denoted as $t_{\text{esc}}$ and calculated using Eq. 2, are presented in Fig. 5a. The general trend of $t_{\text{esc}}$ with cluster size broadly exhibits two distinct stages, each reflecting a different vacancy trapping characteristic. In the initial stage, as cluster size increases, vacancy escape time increases sharply on a logarithmic scale, from approximately $10^{-3}$ seconds to $10^{1}$ seconds. This rapid growth indicates that smaller clusters significantly enhance their vacancy trapping ability with growth. This trend aligns with the observation that the energy basin depth (i.e. the lowest site energy within the cluster) increases linearly with the size for smaller clusters [33].
@@
However, as cluster size exceeds roughly 100 atoms, this trend transitions to a second stage, where further increases in cluster size lead to only marginal changes in $t_{\text{esc}}$. This behavior indicates that the vacancy trapping becomes relatively insensitive to size for large clusters. The underlying reason is that the depth of the PEL basin saturates, converging to the vacancy formation free energy difference between the Al matrix and the solute-rich secondary phase [8, 10, 11]. Beyond this point, continued growth increases only the width rather than the depth of the energy basin. Although the width effect, associated with additional random walk steps that a vacancy must take to reach the cluster boundary, is typically less dominant than the depth effect due to relatively low internal migration barriers, it is not negligible. As the number of internal steps increases, we still observe that the vacancy escape time continues to grow.
@@
Now, our absorbing Markov chain model captures both the depth and the width of the energy basin, incorporating the combined effects of both thermodynamic binding energies and kinetic migration barriers within the cluster. With the vacancy escape times (Fig. 5a), we can more comprehensively characterize the vacancy trapping mechanisms in terms of energetics. We adopt a singularity approximation of the PEL, illustrated in Fig. 5b, where the cluster is simplified as a single site with strong vacancy binding instead of representing the cluster as a complex energy basin of the PEL. This simplification allows us to quantify the trapping ability of the cluster through a single effective binding energy value.
We define the effective cluster energy barrier, $\tilde{E}^{\text{clu}}_{\text{barrier}}$, which governs the escape time via the Arrhenius relationship:
@@
As indicated by Fig. 5b, $\tilde{E}^{\text{clu}}_{\text{barrier}}$ can be larger than $E_{\text{mig}}$ so that a negative value of $\tilde{E}^{\text{clu}}_{\text{bind}}$ indicate that a vacancy is effectively trapped by a solute cluster. Fig. 5c plots $\tilde{E}^{\text{clu}}_{\text{bind}}$ as a function of the cluster size, with dashed lines indicating numerical fits. These results quantitatively confirm that for large clusters, the $\tilde{E}^{\text{clu}}_{\text{bind}}$ increases rapidly and approaches a size-independent limit due to the saturation of binding strength within the interior. Nevertheless, even in this saturation stage, the effective binding energy continues to capture the impact of the random walk behavior of the vacancies within the cluster.
@@
In multicomponent alloys, the vacancy concentration is influenced not only by heat treatment conditions but also by the chemical composition and spatial distribution of solute atoms. These solute atoms can aggregate to form clusters that effectively bind vacancies, thereby altering the concentration of mobile vacancies in the lattice. The effective binding energy to a given solute cluster, denoted as $\tilde{E}^{\text{clu}}_{\text{bind}}$, derived from the absorbing Markov chain model, offers a rigorous and quantitative framework for evaluating the impact of various types and sizes of solute clusters on excess vacancy concentrations. Yang et al. [51] reviewed prior studies [8, 9, 10, 52, 53] and summarized the equilibrium vacancy concentration as a function of temperature. In this framework, the total number of vacancies is categorized into three distinct groups:
@@
The first group, $c^{\text{vac,X}}_{\text{eq}}$, denotes the atomic fraction of vacancies trapped by individual solute atom X. The reference vacancy concentration in pure Al is represented by $c^{\text{vac,Al}}_{\text{eq}}$. Both this quantity and the concentration of isolated solute atoms, $c^{\text{X}}_{1}$, are expressed as atomic fractions. The parameter $z$ represents the coordination number, i.e., the number of nearest-neighbor sites surrounding a solute atom. The term $E_{\text{bind}}^{\text{X}}$ denotes the vacancy binding energy associated with solute atom X.
$c^{\text{vac,X}}_{\text{eq}}=c^{\text{vac,Al}}_{\text{eq}}\left(T\right)\sum_{\text{X}}\left[zc^{\text{X}}_{1}\exp\left(\frac{E_{\text{bind}}^{{\text{X}}}}{k_{B}T}\right)\right],$ (9)
The second group, $c^{\text{vac,clu}}_{\text{eq}}$, is the fraction of vacancies trapped by clusters. The sizes of these clusters are expressed as $N^{\text{clu}}$, which is the number of lattice sites occupied by this cluster. The concentrations of these clusters, $c^{\text{clu}}$, are expressed as an atomic fraction (number of clusters in the whole simulation supercell divided by the number of lattice sites in this supercell). The term $z^{\text{clu}}$ represents the number of neighboring sites which are direct nearest neighbors to any atom belonging to the cluster. These sites provide additional sites for trapping vacancies on the cluster’s surface. $\tilde{E}_{\text{bind}}^{\text{clu}}$ is the effective vacancy binding energy of clusters that can be provided by Eq. 7.
$c^{\text{vac,clu}}_{\text{eq}}=c^{\text{vac,Al}}_{\text{eq}}\left(T\right)\sum_{\text{clu}}\left[\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}\exp\left(\frac{\tilde{E}_{\text{bind}}^{\text{clu}}}{k_{B}T}\right)\right],$ (10)
@@
$c^{\text{vac,res}}_{\text{eq}}=c^{\text{vac,Al}}_{\text{eq}}\left(T\right)\left[1-\sum_{\text{X}}{\left(z+1\right)c^{\text{X}}}-\sum_{\text{clu}}\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}\right],$ (11)
Although Eq. 11 offers a refined estimate of the equilibrium residual vacancy concentration, the vacancy concentration during quenching remains far from equilibrium due to rapid kinetic processes. Recent computational and experimental studies [4, 5] have demonstrated that vacancy supersaturation can persist for several hours to days ($\sim 10^{5}$-$10^{6}$ seconds) in regions far from GB. Therefore, we focus on the interactions between vacancies and both solute atoms and solute clusters, while assuming that the total vacancy concentration remains constant during quenching in large-grained regions. The expression for the dynamic vacancy concentration is given by:
$c^{\text{vac,res}}_{\text{dyn}}=c^{\text{vac}}_{\text{eq}}\left(800K\right)\frac{\sum_{\text{Al}}\exp\left({-\frac{E_{\text{bind}}^{\text{site}}}{k_{B}T}}\right)}{\sum_{\text{all}}\exp\left({-\frac{E_{\text{bind}}^{\text{site}}}{k_{B}T}}\right)},$ (12)
where the numerator and denominator represent the total probability of finding vacancies in all lattice sites of the Al matrix and in all lattice sites of the entire alloy, respectively, where $E_{\text{bind}}^{\text{site}}$ is the vacancy binding energy of each lattice site. It is important to note that the local binding energy of sites in the pure Al matrix is defined to be zero. By substituting the effective cluster binding energy, $\tilde{E}_{\text{bind}}^{\text{clu}}$, for $E_{\text{bind}}^{\text{site}}$ for each site belong to clusters, Eq. 12 can be reformulated as:
$c^{\text{vac,res}}_{\text{dyn}}=c^{\text{vac}}_{\text{eq}}\left(800K\right)\frac{1-\sum_{\text{clu}}\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}}{1-\sum_{\text{clu}}\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}+\sum_{\text{clu}}\left[\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}\exp\left(\frac{\tilde{E}_{\text{bind}}^{\text{clu}}}{k_{B}T}\right)\right]}.$ (13)
@@
Fig. 6a displays the temperature evolution profiles under three different cooling conditions: instantaneous quenching, fast cooling, and slow cooling (see Supplementary Note 4 for cluster snapshots). Fig. 6b presents the corresponding evolution of residual vacancy concentrations, including both the equilibrium concentration $c^{\text{vac,res}}_{\text{eq}}$ (from Eq. 11) and the dynamic concentration $c^{\text{vac,res}}_{\text{dyn}}$ (from Eq. 13) under these three conditions. In all cases, $c^{\text{vac,res}}_{\text{dyn}}$ remains higher than $c^{\text{vac,res}}_{\text{eq}}$ because equilibrium has not yet been reached. For the instantaneous quenching case, solute clusters cannot grow rapidly, leading to fewer large clusters capable of effectively trapping vacancies. As a result, a higher $c^{\text{vac,res}}_{\text{dyn}}$ is observed in the Al matrix. Over time, as clustering progresses and larger clusters form, $c^{\text{vac,res}}_{\text{dyn}}$ gradually decreases. In contrast, under more realistic cooling conditions (fast and slow cooling), stronger vacancy-cluster interactions are observed due to the formation of larger clusters during the cooling process. These clusters absorb more vacancies through strong binding interactions, resulting in lower $c^{\text{vac,res}}_{\text{dyn}}$ at a given time. Overall, $c^{\text{vac,res}}_{\text{dyn}}$ tends to be higher for faster cooling rates than for slower ones, as faster cooling suppresses the formation of additional large clusters that trap vacancies more effectively. Interestingly, this trend is opposite to that observed for $c^{\text{vac,res}}_{\text{eq}}$, where faster cooling corresponds to a lower temperature at a given time, thereby yielding lower equilibrium vacancy concentrations. This opposite trend can be rationalized by noting that faster cooling rates lead to a larger deviation between the dynamic and equilibrium vacancy concentrations, i.e., a greater gap between $c^{\text{vac,res}}_{\text{dyn}}$ and $c^{\text{vac,clu}}_{\text{eq}}$, due to the system being driven further from equilibrium by rapid kinetics.
Figure 6: Evolution of residual vacancy concentration during cooling, estimated from KMC simulations under three different cooling rates. (a) Temperature profiles corresponding to the three cooling conditions used in the KMC simulations. (b) Comparison of equilibrium residual vacancy concentration ($c^{\text{vac,res}}_{\text{eq}}$ defined in Eq. 11, dashed lines) and residual vacancy concentration ($c^{\text{vac,res}}_{\text{dyn}}$ defined in Eq. 13, solid lines) over time for instantaneous, fast, and slow cooling scenarios.
@@
Predictions of $c^{\text{vac,res}}_{\text{dyn}}$, the concentration of mobile vacancies available for solute diffusion in the Al matrix, provide a solid foundation for studying the nucleation and growth kinetics of solute clusters in Al alloys. The diffusivity of solute species, $D^{\text{X}}$, is directly influenced by vacancy concentration and migration barriers, and follows the relationship $D^{\text{X}}\propto c^{\text{vac,res}}_{\text{dyn}}\exp\left(-{E_{\text{mig}}^{\text{X}}}/{k_{B}T}\right)$. This diffusivity governs both the nucleation rate, $dN/dt\propto D^{\text{X}}\exp(-G^{*}/{k_{B}T})$, and the growth rate, $dR/dt\propto D^{\text{X}}/R$, which are critical for describing precipitation kinetics. The KMC results shown in Fig. 1 demonstrate that traditional KMC simulations encounter challenges in efficiently simulating cluster growth beyond the early stages. These limitations arise due to strong vacancy-cluster interactions and the finite number of vacancies in a typical simulation supercell. To overcome these constraints and to study clustering behavior over longer timescales relevant to natural aging, CD simulations were employed. By incorporating the accurately determined vacancy supersaturation obtained from the absorbing Markov chain model, the CD framework (see Methods section for details) bridges atomistic-scale vacancy interactions with macroscopic precipitation processes, enabling robust modeling of cluster size distribution evolution over extended time periods.

### v2

@@ Results
where $E^{\text{site}}_{\text{bind},i}$ is the energy of vacancy on node $i$. The denominator sums contributions from all $n_{\text{tr}}$ transient states. The element, $\tau_{i}$, in the vector $\vec{\tau}$ represents the expected escape one-step time for the particular site. For each site $i$, $\tau_{i}$ accounts for all possible transitions to neighboring states. Migration barriers $E^{\text{bond}}_{\text{mig},ij}$, and attempt frequencies $\nu$ are incorporated to estimate escape times from each transient state:
$\tau_{i}=\left(\nu\sum_{j=1}^{z}{\exp\left(-\frac{E^{\text{bond}}_{\text{mig},ij}}{k_{B}T}\right)}\right)^{-1},$ (5)
where $\nu$ is set at $10^{13}$ Hz, as a typical atomic vibration frequency value in a crystal.
@@
Figure 4: KMC simulations for the verification of the absorbing Markov chain model described by Eq. 2. (a) Illustration of the KMC strategy used to determine the vacancy escape event from a solute cluster. Red open circles represent vacancies, while blue, green, and orange circles indicate Al, Mg, and Zn atoms, respectively. White stars mark the first nearest-neighbor shell surrounding the cluster, which is used to define the escape condition. (b) Comparison of vacancy escape times, $t_{\text{esc}}$, obtained from the analytical solution of Eq. 2 and from direct KMC simulations for solute clusters generated by SA simulations with varying cluster sizes, $N^{\text{clu}}$.
@@
For solute clusters in multicomponent systems, such as Mg-Zn-rich clusters in Al-Mg-Zn alloys, we recognize that local atomic rearrangements by vacancy migration may change the actual PEL. Despite these structural changes, we make a simplified-PEL assumption. Specifically, we assume that the PEL experienced by a migrating vacancy can be approximated by the static ground-state structures of the solute cluster (obtained by SA simulations). This simplification is reasonable for solute clusters with lower vacancy migration barriers within the clusters (as shown in Fig. 2) at lower temperatures, where low-energy cluster structures are favorable by considering both thermodynamic and kinetic factors. By making this assumption, we effectively reduce the complexity of the PEL, enabling a computationally efficient absorbing Markov chain model to describe vacancy dynamics inside the clusters. This simplification maintains analytical tractability and computational efficiency, effectively capturing essential vacancy behavior without explicitly modeling structural rearrangements.
@@
We applied KMC simulations to validate the absorbing Markov chain model and test the assumption of a simplified PEL. KMC and Markov chains share the same memorylessness property, which assumes that the next state depends only on the current state and not on past states. KMC calculates transition rates and selects events probabilistically, updating the system state and time step by step. Unlike the Markov model, KMC inherently accounts for structural rearrangements and complex PELs by updating the lattice occupations after each vacancy migration event.
@@
Fig. 4 presents the verification of our absorbing Markov chain model based on the simplified PEL of vacancy migrations inside solute clusters, which are generated and optimized from the SA simulations (Supplementary Note 3 describes the same verification processes applied to solute clusters directly extracted from KMC simulations). Each cluster contains one vacancy and has a fixed Zn/Mg ratio of 2 due to the SA simulation optimization. Then, we performed KMC simulations on these clusters, focusing on determining the real escape times required for a vacancy to leave the cluster. In KMC simulations, the vacancies are tracked as they migrate through the cluster until they reach a site outside of the first-nearest-neighbor shell, where they are considered to have escaped from this cluster. This stopping criterion, illustrated in Fig. 4a, mirrors the definition used in the absorbing Markov chain model, providing a consistent basis for comparison between the two approaches. The KMC simulations were conducted at an elevated temperature of 450 K. This higher temperature facilitates fast transitions, enabling vacancies to overcome energy barriers and escape the energy basin more efficiently compared to simulations performed at room temperature. This higher temperature ensures that statistically meaningful results can be obtained within a feasible computational time. Fig. 4b compares the escape times predicted by the absorbing Markov chain model (Eq. 2) with those obtained from the KMC simulations for clusters with varying cluster sizes, $N^{\text{clu}}$, which is the number of atoms that belong to this cluster. Here, results from absorbing Markov chain and KMC validation exhibit strong agreement, demonstrating that the simplified PEL approximation employed in the absorbing Markov model effectively captures the essential features of vacancy dynamics.
@@
Figure 5: Analysis for vacancy trapping ability of solute clusters using the Markov chain model. (a) Vacancy escape time $t_{\text{esc}}$ calculated by Eq. 2 at 300 K of solute clusters in terms of cluster size $N^{\text{clu}}$. Multiple cluster configurations with the same $N^{\text{clu}}$ value were analyzed to statistically represent the trapping behavior. (b) A conceptual figure to explain the single-energy-basin approximation for a solute cluster. (c) Effective vacancy binding energy of a cluster $\tilde{E}^{\text{clu}}_{\text{bind}}$ defined in Eq. 7 in terms of cluster size $N^{\text{clu}}$. The dashed line represents a numerical fit.
@@
To investigate the influence of vacancy-cluster interactions on long-time solute clustering kinetics during natural aging, we calculated vacancy escape times, $t_{\text{esc}}$, at 300 K using Eq. 2 for solute clusters generated by the SA simulations with varying $N^{\text{clu}}$. The results, shown in Fig. 5a, reveal two distinct regimes in the dependence of $t_{\text{esc}}$ on $N^{\text{clu}}$, each reflecting different vacancy trapping behaviors. In the initial regime, as the cluster size increases, the vacancy escape time rises sharply on a logarithmic scale, from approximately $10^{-3}$ s to $10^{1}$ s. This rapid growth indicates that small clusters strongly enhance their vacancy-trapping capability as they grow. This trend is consistent with previous studies showing that the depth of the energy basin, associated with the lowest site binding energy within the cluster, increases approximately linearly with cluster size for small clusters [38].
@@
However, once the cluster size exceeds approximately 100 atoms, the trend transitions to a second regime, where further increases in $N^{\text{clu}}$ yield only modest changes in $t_{\text{esc}}$. In this regime, the vacancy trapping becomes relatively insensitive to cluster size. This behavior arises because the depth of the PEL basin saturates, approaching the vacancy formation energy difference between the Al matrix and the solute-rich precipitate phase [13, 15, 16]. Beyond this point, additional cluster growth primarily increases the basin width rather than its depth. Although this widening effect, associated with the increased number of random walk steps a vacancy must take to escape the cluster, is generally weaker than the depth effect due to relatively low internal migration barriers, it is still non-negligible. As a result, vacancy escape times continue to increase gradually with growing cluster size.
@@
Now, our absorbing Markov chain model captures both the depth and the width of the energy basin, incorporating the combined effects of both thermodynamic binding energies and kinetic migration barriers within the cluster. With the vacancy escape times (Fig. 5a), we can more comprehensively characterize the vacancy trapping mechanisms in terms of energetics. We adopt a single-energy-basin approximation of the PEL, illustrated in Fig. 5b, where the cluster is simplified as a single site with strong vacancy binding instead of representing the cluster as a complex energy basin of the PEL. This simplification allows us to quantify the trapping ability of the cluster through a single effective binding energy value. Thus, we define the effective cluster energy barrier, $\tilde{E}^{\text{clu}}_{\text{barrier}}$, which governs the escape time via the Arrhenius relationship:
@@
As indicated by Fig. 5b, $\tilde{E}^{\text{clu}}_{\text{barrier}}$ can be larger than $E_{\text{mig}}$ so that a negative value of $\tilde{E}^{\text{clu}}_{\text{bind}}$ indicate that a vacancy is effectively trapped by a solute cluster. Fig. 5c plots $\tilde{E}^{\text{clu}}_{\text{bind}}$ as a function of the cluster size $N^{\text{clu}}$. Due to the stochastic nature of the Monte Carlo-based SA methods and the inherent complexity of the cluster PELs, it is challenging for individual simulations to consistently identify the lowest-energy configurations. Consequently, multiple configurations were analyzed to statistically represent the trapping behavior. A numerical fit (the dashed line in Fig. 5b) is applied to capture the representative relationship between effective binding energies $\tilde{E}^{\text{clu}}_{\text{bind}}$ and cluster sizes, accounting for variability among different cluster configurations. These results quantitatively confirm that for large clusters, the $\tilde{E}^{\text{clu}}_{\text{bind}}$ increases rapidly and approaches a size-independent limit due to the saturation of binding strength within the interior. Nevertheless, even in this saturation stage, the effective binding energy continues to capture the impact of the random walk behavior of the vacancies within the cluster.
@@
In multicomponent alloys, the long-time solute clustering kinetics during quenching and natural aging are strongly governed by the concentration of residual vacancies in the matrix that remain available for solute diffusion. These residual vacancy concentrations are significantly influenced by interactions between vacancies and solute atoms (or solute clusters), which can effectively trap vacancies and thus reduce the number of mobile vacancies in the matrix. The effective vacancy binding energy to a given solute cluster, denoted as $\tilde{E}^{\text{clu}}_{\text{bind}}$ in Eq. 7, provides a rigorous and quantitative framework to evaluate the influence of various solute cluster types and sizes on residual vacancy concentrations as described in the following analyses.
Based on previous studies [13, 14, 15, 60, 61, 62], we classify the equilibrium vacancy concentration as a function of temperature, $c^{\text{vac}}_{\text{eq}}\left(T\right)$, into three distinct categories, which are described as follows:
@@
The first group, $c^{\text{vac,X}}_{\text{eq}}$, denotes the atomic fraction of vacancies trapped by individual solute atom X isolated in the Al matrix as follows:
$c^{\text{vac,X}}_{\text{eq}}\left(T\right)=c^{\text{vac,Al}}_{\text{eq}}\left(T\right)\sum_{\text{X}}\left[zc^{\text{X}}_{1}\exp\left(-\frac{E_{\text{bind}}^{{\text{X}}}}{k_{B}T}\right)\right],$ (9)
The reference vacancy concentration in pure Al is represented by $c^{\text{vac,Al}}_{\text{eq}}$. Both this quantity and the concentration of isolated solute atoms, $c^{\text{X}}_{1}$, are expressed as atomic fractions. The parameter $z$ represents the coordination number, i.e., the number of nearest-neighbor sites surrounding a solute atom. The term $E_{\text{bind}}^{\text{X}}$ denotes the vacancy binding energy associated with this individual solute atom X, where a negative value signifies attractive binding.
The second group, $c^{\text{vac,clu}}_{\text{eq}}$, is the fraction of vacancies trapped by solute clusters as follows:
$c^{\text{vac,clu}}_{\text{eq}}\left(T\right)=c^{\text{vac,Al}}_{\text{eq}}\left(T\right)\sum_{\text{clu}}\left[\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}\exp\left(-\frac{\tilde{E}_{\text{bind}}^{\text{clu}}}{k_{B}T}\right)\right],$ (10)
The sizes of these clusters are expressed as $N^{\text{clu}}$, which is the number of lattice sites occupied by this cluster. The concentrations of these clusters, $c^{\text{clu}}$, are expressed as an atomic fraction (the number of clusters in the whole simulation supercell divided by the number of lattice sites in this supercell). The term $z^{\text{clu}}$ represents the number of neighboring sites that are direct nearest neighbors to any atom belonging to the cluster. These sites provide additional sites for trapping vacancies at the cluster-matrix interfaces. $\tilde{E}_{\text{bind}}^{\text{clu}}$ is the effective vacancy binding energy of clusters calculated by Eq. 7, with negative values indicating attractive binding.
@@
$c^{\text{vac,res}}_{\text{eq}}\left(T\right)=c^{\text{vac,Al}}_{\text{eq}}\left(T\right)\left[1-\sum_{\text{X}}{\left(z+1\right)c^{\text{X}}_{1}}-\sum_{\text{clu}}\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}\right],$ (11)
Although Eq. 11 offers a refined estimate of the equilibrium residual vacancy concentration in the Al matrix, the vacancy concentration during quenching remains far from equilibrium due to rapid kinetic processes. Recent computational and experimental studies [4, 5] have demonstrated that vacancy supersaturation can persist for several hours to days ($\sim 10^{5}$-$10^{6}$ seconds) in regions far from GB. Therefore, we focus on the interactions between vacancies and both solute atoms and solute clusters, while assuming that the total vacancy concentration remains constant during quenching in large-grained regions. The expression for the dynamic vacancy concentration is given by:
$c^{\text{vac,res}}_{\text{dyn}}\left(T\right)=c^{\text{vac}}_{\text{eq}}\left(800K\right)\frac{\sum_{\text{Al}}\exp\left({-\frac{E_{\text{bind}}^{\text{site}}}{k_{B}T}}\right)}{\sum_{\text{all}}\exp\left({-\frac{E_{\text{bind}}^{\text{site}}}{k_{B}T}}\right)},$ (12)
where the numerator and denominator represent the total probability of finding vacancies in all lattice sites of the Al matrix and in all lattice sites of the entire alloy, respectively, with $E_{\text{bind}}^{\text{site}}$ as the vacancy binding energy at the corresponding site. It is important to note that the local binding energy of sites in the pure Al matrix is defined to be zero. By substituting the effective cluster binding energy, $\tilde{E}_{\text{bind}}^{\text{clu}}$, for $E_{\text{bind}}^{\text{site}}$ for each site belong to clusters, Eq. 12 can be reformulated as:
$c^{\text{vac,res}}_{\text{dyn}}\left(T\right)=\frac{c^{\text{vac}}_{\text{eq}}\left(800K\right)\left(1-\sum_{\text{X}}{\left(z+1\right)c^{\text{X}}_{1}}-\sum_{\text{clu}}\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}\right)}{1-\sum_{\text{X}}{\left[\left(z+1\right)c^{\text{X}}_{1}-zc^{\text{X}}_{1}\exp\left(-\frac{E_{\text{bind}}^{{\text{X}}}}{k_{B}T}\right)\right]}-\sum_{\text{clu}}{\left[\left(z^{\text{clu}}+N^{\text{clu}}\right)c^{\text{clu}}\left(1-\exp\left(-\frac{\tilde{E}_{\text{bind}}^{\text{clu}}}{k_{B}T}\right)\right)\right]}}.$ (13)
@@
Fig. 6a displays the temperature evolution profiles under three different cooling conditions: instantaneous quenching, fast cooling, and slow cooling (see Supplementary Note 4 for cluster snapshots). Here, cluster structures and effective vacancy-cluster interaction energies were calculated from KMC simulations and the Markov chain model described above. Fig. 6b presents the corresponding evolution of residual vacancy concentrations in the Al matrix, including both the equilibrium concentration $c^{\text{vac,res}}_{\text{eq}}$ (from Eq. 11) and the dynamic concentration $c^{\text{vac,res}}_{\text{dyn}}$ (from Eq. 13) under these three conditions. In all cases, $c^{\text{vac,res}}_{\text{dyn}}$ remains higher than $c^{\text{vac,res}}_{\text{eq}}$ because equilibrium has not yet been reached. For the instantaneous quenching case, solute clusters cannot grow rapidly, leading to fewer large clusters capable of effectively trapping vacancies. As a result, a higher $c^{\text{vac,res}}_{\text{dyn}}$ is observed in the Al matrix. Over time, as clustering progresses and larger clusters form, $c^{\text{vac,res}}_{\text{dyn}}$ gradually decreases. In contrast, under more realistic cooling conditions (fast and slow cooling), stronger vacancy-cluster interactions are observed due to the formation of larger clusters during the cooling process. These clusters absorb more vacancies through strong binding interactions, resulting in lower $c^{\text{vac,res}}_{\text{dyn}}$ at a given time. Overall, $c^{\text{vac,res}}_{\text{dyn}}$ tends to be higher for faster cooling rates than for slower ones, as faster cooling suppresses the formation of additional large clusters that trap vacancies more effectively. Interestingly, this trend is opposite to that observed for $c^{\text{vac,res}}_{\text{eq}}$, where faster cooling corresponds to a lower temperature at a given time, thereby yielding lower equilibrium vacancy concentrations. This opposite trend can be rationalized by noting that faster cooling rates lead to a larger deviation between the dynamic and equilibrium vacancy concentrations, i.e., a greater gap between $c^{\text{vac,res}}_{\text{dyn}}$ and $c^{\text{vac,clu}}_{\text{eq}}$, due to the system being driven further from equilibrium by rapid kinetics.
Figure 6: Evolution of residual vacancy concentration during quenching under three different cooling rates, where the corresponding cluster structures and effective vacancy-cluster interaction energies estimated from KMC simulations and the Markov chain model. (a) Temperature profiles for the three cooling conditions used in the KMC simulations. (b) Time evolution of the equilibrium residual vacancy concentration ($c^{\text{vac,res}}_{\text{eq}}$, defined in Eq. 11, shown as dashed lines) and the dynamic residual vacancy concentration ($c^{\text{vac,res}}_{\text{dyn}}$, defined in Eq. 13, shown as solid lines) for the instantaneous, fast, and slow cooling scenarios presented in (a).
@@
Predictions of $c^{\text{vac,res}}_{\text{dyn}}$, the concentration of mobile vacancies available for solute diffusion in the Al matrix, provide a solid foundation for studying the long-time solute clustering kinetics in Al alloys. The diffusivity of solute species, $D^{\text{X}}$, is directly influenced by vacancy concentration and migration barriers, and follows the relationship $D^{\text{X}}\propto c^{\text{vac,res}}_{\text{dyn}}\exp\left(-{E_{\text{mig}}^{\text{X}}}/{k_{B}T}\right)$. This diffusivity governs both the nucleation and growth rates, which are critical for describing precipitation kinetics. The KMC results shown in Fig. 1 demonstrate that traditional KMC simulations encounter challenges in efficiently simulating cluster growth beyond the early stages. These limitations arise due to strong vacancy-cluster interactions and the finite number of vacancies in a typical simulation supercell. To overcome these constraints and to study clustering behavior over longer timescales during natural aging, CD simulations were employed. By incorporating the accurate vacancy supersaturation obtained from the absorbing Markov chain model, the CD framework (see Methods section for details) bridges atomistic-scale vacancy-cluster interaction energetics with macroscopic precipitation processes, enabling robust modeling of cluster size distribution evolution over extended time periods.